# Cybersecurity Challenge Analysis of Work-from-Anywhere (WFA) and Recommendations guided by a User Study Authors’ draft for soliciting feedback, September 11, 2024

### v1

@@ Abstract
Many organizations were forced to quickly transition to the work-from-anywhere (WFA) model as a necessity to continue with their operations and remain in business despite the restrictions imposed during the COVID-19 pandemic. Many decisions were made in a rush, and cybersecurity decency tools were not in place to support this transition. In this paper, we first attempt to uncover some challenges and implications related to the cybersecurity of the WFA model. Secondly, we conducted an online user study to investigate the readiness and cybersecurity awareness of employers and their employees who shifted to work remotely from anywhere. The user study questionnaire addressed different resilience perspectives of individuals and organizations. The collected data includes responses from remotely working employees of different organizational types: universities, government, private, and non-profit organizations.
Despite the importance of security training and guidelines, it was surprising that many participants had not received them. A robust communication strategy is necessary to ensure that employees are informed and updated on security incidents that the organization encounters. Additionally, there is an increased need to pay attention to the security-related attributes of employees, such as their behavior, awareness, and compliance. Finally, we outlined best practice recommendations and mitigation tips guided by the study results to help individuals and organizations resist cybercrime and fraud and mitigate WFA-related cybersecurity risks.
@@
The work-from-anywhere (WFA) model has become increasingly popular in recent years, allowing employees to work remotely and have more flexible schedules. A Gartner survey [1] reveals that more than of organizations will support a certain population of their workforce to continue working from home or anywhere. However, this model introduces new risks and challenges that transform the cybersecurity landscape. Attackers may exploit the vulnerabilities that arise from employee behavior and employer policies [2]. Remote employees may connect to unsecured public WiFi networks, use shadow IT [3], and personal devices to access corporate resources or store sensitive data on unsecured devices, increasing the risk of data breaches and cyber-attacks. Phishing attacks, malware infections, and insider threats can also be more challenging to detect and prevent when employees work remotely. Additionally, compliance with regulatory requirements can be challenging when employees work remotely. By understanding these challenges and implementing appropriate cybersecurity measures, organizations can ensure that the WFA model is a safe and secure environment for their employees and their data.
In light of the above discussion, the contribution of this paper is threefold:
@@
The remainder of this paper is structured as follows. Section II reviews the related work on the topic under study. Section III presents the main components of the WFA model and its common challenges and implications related to cybersecurity. Section IV presents the design methodology, demographic, and validity analysis of the user study. The data collected are analyzed and discussed in Section V. In Section VI, the key findings and limitations of this study are given. Section VII presents recommendations for effective cyberattack risk reduction practices to enhance the WFA-related cybersecurity culture. Finally, Section VIII concludes the paper.
@@ IV-A User-Study Designing and Testing
The user study conducted in this work attempts to address the following questions in the context of WFA: Q1: What are the security precautions employers are practicing?, Q2: What is the level of security awareness among employees regarding security threats and best practices? Q3: To what extent can employee behavior affect the overall security of the organization?.
To obtain answers to our research questions, we conducted an anonymous online user study approved by CUREB111https://carleton.ca/researchethics/cureb-b/, CUREB-B Clearance#117292. The study targeted employees who work remotely, fully or partially, during the pandemic. Participants were asked to complete a questionnaire of questions divided into four parts, designed to minimize cognitive load and allow participants to fully consider the topics presented [30]. The first part of the questionnaire collected demographic information, working environment, work experience, IT knowledge, work sector, and the size of the organization’s manpower. The second part focused on organizations’ preparedness to shift to the WFA scheme and their procedures to defend against detected attacks and increase employee security awareness. The third part aimed to understand common security-related behaviors of employees while working remotely. The final part sought to determine employees’ awareness of common cyberattacks and vulnerabilities. We follow the recommended three-stage process by Dillman [31] to pre-test the questionnaire. First, the authors’ team discussed the clarity of the questionnaire and the motivation to identify unclear questions, ambiguous instructions, or other issues before public dissemination. Colleagues and experts in the field then reviewed the questionnaire to uncover potential misunderstandings or unexpected outcomes. Finally, we performed pilot tests with eight graduate-level colleagues to identify any flaws in the questionnaire and determine its appropriate length. The user study was finalized after incorporating feedback from the previous stages. We used Qualtrics [32] as an online survey tool for this study.
@@ V-C5 Lack of communication between employers and employees about cyberattacks
The study found that there is a miscommunication between employers and employees regarding cyberattacks. A descriptive analysis was conducted to determine whether employers usually inform their employees about attacks that occurred during the pandemic. The results show that of the participants from all types of organizations do not receive any notification. Figure 14 illustrates the number of participants who receive information and those who do not. Interestingly, of government employees do not get informed about attacks, representing the highest percentage among participants in different organizations. It is important to note that although there is a possibility that organizations do not inform their employees due to the absence of attacks, it is unlikely that all organizations have remained unaffected by any attacks.
@@ VII Best practice recommendations
VII-1 Enhancing help desk
@@
VII-2 Importance of Antivirus Software
@@
VII-3 Implementing secure network segmentation for work activities
@@
VII-4 Promoting secure behaviors and limit risky practices
@@
VII-5 Avoiding public WiFi if possible
@@ VII-6 Using VPN
VII-7 Using personal hotspots
@@ VII-8 Implementing MFA
VII-9 Establishing clear and updated security policies
@@
VII-10 Providing comprehensive security awareness training
@@
VII-11 Developing an incident response plan
@@
VII-12 Establishing better communication channels with employees
@@
VII-13 Encouraging reporting of suspicious activities
@@ VIII Conclusion
Given the effects of the increase in cyberattacks internationally during COVID-19, understanding the rapid changes in the working environment on IT and information security is essential. Understanding the difficulties posed by such significant changes in security management lays the path for improved crisis readiness and a secure transition to remote work. In this paper, we analyzed WFA-related cybersecurity challenges, conducted a user study to evaluate the preparedness and cybersecurity awareness of organizations and individuals to shift to work remotely from anywhere, and outlined best practice recommendations to reduce cybersecurity risks.

### v2

@@ Abstract
Many organizations were forced to quickly transition to the work from anywhere (WFA) model as a necessity to continue with their operations and remain in business despite the restrictions imposed during the COVID-19 pandemic. Many decisions were made in a rush, and cybersecurity decency tools were not in place to support this transition. In this paper, we first attempt to uncover some challenges and implications related to the cybersecurity of the WFA model. Secondly, we conducted an online user study to investigate the readiness and cybersecurity awareness of employers and their employees who shifted to work remotely from anywhere. The user study questionnaire addressed different resilience perspectives of individuals and organizations. The collected data includes responses from remotely working employees of different organizational types: universities, government, private, and non-profit organizations. Despite the importance of security training and guidelines, it was surprising that many participants had not received them. A robust communication strategy is necessary to ensure that employees are informed and updated on security incidents that the organization encounters. Additionally, there is an increased need to pay attention to the security-related attributes of employees, such as their behavior, awareness, and compliance. Finally, we outlined best practice recommendations and mitigation tips guided by the study results to help individuals and organizations resist cybercrime and fraud and mitigate WFA-related cybersecurity risks.
@@
The work-from-anywhere (WFA) model has become increasingly popular in recent years, allowing employees to work remotely and have more flexible schedules. A Gartner survey [1] reveals that more than of organizations will support a certain population of their workforce to continue working from home or anywhere. However, this model introduces new risks and challenges that transform the cybersecurity landscape. Attackers may exploit the vulnerabilities that arise from employee behavior and employer policies [2]. Remote employees may connect to unsecured public WiFi networks, use shadow IT [3], and personal devices to access corporate resources or store sensitive data on unsecured devices, increasing the risk of data breaches and cyber-attacks. Phishing attacks, malware infections, and insider threats can also be more challenging to detect and prevent when employees work remotely. Additionally, compliance with regulatory requirements can be challenging when employees work remotely. By understanding these challenges and implementing appropriate cybersecurity measures, organizations can ensure that the WFA model is a safe and secure environment for their employees and their data. However, despite these increasing risks, there is a well-documented discrepancy between individuals’ privacy concerns and their actual behaviour, often referred to as the privacy paradox as reviewed by Gerber et al. [5]. This can be explained by several factors, including behavioural biases (such as the intention-behaviour gap), a lack of awareness of privacy risks, and trade-offs between privacy and other perceived benefits, such as convenience. In the context of WFA, this paradox is particularly relevant, as remote employees may express concerns about cybersecurity but still engage in risky behaviours, such as using public networks without using VPN due to convenience or lack of perceived immediate threat. In light of the above discussion, the contribution of this paper is threefold:
@@
The remainder of this paper is structured as follows. Section II reviews the related work. Section III presents the WFA model components and its cybersecurity challenges and implications. Section IV presents the design methodology, demographic, and validity analysis of the user study. The data collected are analyzed in Section V. In Section VI, the key findings and limitations of this study are given. Section VII presents best practice recommendations for effective cyberattack risk reduction. Finally, Section VIII concludes the paper.
@@ IV-A User-Study Designing and Testing
The user study conducted in this work attempts to address the following questions in the context of WFA: Q1: What are the security precautions employers are practicing?, Q2: What is the level of security awareness among employees regarding security threats and best practices? Q3: To what extent can employee behavior affect the overall security of the organization?. To obtain answers to our research questions, we conducted an anonymous online user study approved by CUREB111https://carleton.ca/researchethics/cureb-b/, CUREB-B Clearance#117292. The study targeted employees who work remotely, fully or partially, during the pandemic. Participants were asked to complete a questionnaire of questions divided into four parts, designed to minimize cognitive load and allow participants to fully consider the topics presented [31]. The first part of the questionnaire collected demographic information, working environment, work experience, IT knowledge, work sector, and the size of the organization’s manpower. The second part focused on organizations’ preparedness to shift to the WFA scheme and their procedures to defend against detected attacks and increase employee security awareness. The third part aimed to understand common security-related behaviors of employees while working remotely. The final part sought to determine employees’ awareness of common cyberattacks and vulnerabilities. We follow the recommended three-stage process by Dillman [32] to pre-test the questionnaire. First, the authors’ team discussed the clarity of the questionnaire and the motivation to identify unclear questions, ambiguous instructions, or other issues before public dissemination. Colleagues and experts in the field then reviewed the questionnaire to uncover potential misunderstandings or unexpected outcomes. Finally, we performed pilot tests with eight graduate-level colleagues to identify any flaws in the questionnaire and determine its appropriate length. The user study was finalized after incorporating feedback from the previous stages. We used Qualtrics [33] as an online survey tool for this study.
@@ IV-D2 Scale reliability of the employees’ satisfaction
IV-E Theoretical Framework
The findings of this study align with well-established theoretical frameworks in cybersecurity and human-computer interaction (HCI). Our analysis can be viewed in the context of the following frameworks: 1) The Protection Motivation Theory (PMT) [38], which explains how individuals adopt protective behaviours when confronted with potential threats. This theory provides insight into how employees perceive cybersecurity risks and take protective actions such as avoiding public Wi-Fi. 2)The Technology Acceptance Model (TAM) [39], which explores how users accept and utilize technology. In the context of WFA, this model helps explain the adoption of security tools provided by employers, such as MFA and VPNs.
@@ V-C5 Lack of communication between employers and employees about cyberattacks
The study found that there is a miscommunication between employers and employees regarding cyberattacks. A descriptive analysis was conducted to determine whether employers generally inform their employees about attacks that occurred during the pandemic. The results show that of the participants from all types of organizations do not receive any notification. Figure 14 illustrates the number of participants who receive information and those who do not. Interestingly, of government employees do not get informed about attacks, representing the highest percentage among participants in different organizations. It is important to note that although there is a possibility that organizations do not inform their employees due to the absence of attacks, it is unlikely that all organizations have remained unaffected by any attacks.
@@ VII-A Recommendations Based on User Study Findings
VII-A1 Enhancing help desk
@@
VII-A2 Importance of Antivirus Software
@@
VII-A3 Implementing secure network segmentation for work activities
@@
VII-A4 Promoting secure behaviors and limit risky practices
@@
VII-A5 Avoiding public WiFi if possible
@@ VII-A6 Using VPN
VII-A7 Using personal hotspots
@@ VII-A8 Implementing MFA
VII-A9 Establishing clear and updated security policies
@@
VII-A10 Providing comprehensive security awareness training
@@
VII-A11 Developing an incident response plan
@@
VII-A12 Establishing better communication channels with employees
@@
VII-A13 Encouraging reporting of suspicious activities
@@
VII-B Findings Alignment with NIST Cybersecurity Framework
While some of these practices may appear basic, our study reveals that many organizations do not consistently implement the most fundamental cybersecurity measures. We compare these recommendations with established NIST security framework to strengthen these recommendations.
1.
Promote secure behavior through training and awareness programs: Our study found gaps in security awareness and behaviour among remote workers, with 40% of the participants reporting they had not received formal security training. By encouraging a security-conscious culture, organizations can enhance employees’ ability to identify and respond to cybersecurity threats. Additionally, this recommendation is aligned with the NIST Cybersecurity Framework’s Protect function (PR.AT), which stresses the need for training to ensure that personnel know their cybersecurity roles and responsibilities.
2.
Implement MFA for remote access: While 70% of organizations in our study had adopted MFA, a portion of respondents reported it was not consistently enforced. Our best practice recommendation is closely aligned with the NIST Cybersecurity Framework’s Protect function (PR.AC), which advocates using MFA to ensure that only authorized individuals can access organizational systems.
3.
Use antivirus software and ensure timely updates: While antivirus software may seem basic, our study found that many participants use it. Antivirus software is a fundamental element of the NIST Cybersecurity Framework’s Protect function (PR.DS), which includes protective technologies such as malware defences to ensure data security.
4.
Ensure regular communication about cybersecurity threats and incidents: Our findings revealed a concerning communication gap, with 79% of respondents reporting they were not informed about cyberattacks or security breaches experienced by their organization. This lack of transparency can hinder an organization’s efforts to encourage a proactive security culture. In comparison, the NIST Cybersecurity Framework’s Respond function (RS.CO) standard both emphasize the importance of clear communication protocols during and after security incidents, ensuring that all relevant stakeholders are informed and prepared to respond to future threats.
@@ VIII Conclusion
This study investigates the cybersecurity challenges of the Work-from-Anywhere (WFA) model, using a structured user study to assess both organizational preparedness and employee behaviors in remote work environments. The findings reveal several critical issues, including a need for more security training for remote workers, communication gaps about cyberattacks, and varying levels of compliance with security protocols across different sectors. Our findings align with the literature on cybersecurity culture, showing how organizational communication and policy dissemination impact employee compliance. The observed communication gaps about cyberattacks highlight the need for a more proactive approach to cybersecurity awareness, particularly in decentralized work environments. Additionally, our recommendations are aligned with the widely accepted NIST Cybersecurity Framework, ensuring that the proposed practices adhere to recognized standards. By contextualizing our findings within NIST’s guidelines, we provide actionable insights that organizations can implement to improve their cybersecurity posture effectively. This alignment enhances the practical relevance of our study and offers a structured pathway for organizations seeking to strengthen their security policies. These theoretical and practical contributions offer a solid foundation for future research exploring the intersection of cybersecurity and remote work while providing actionable guidance for organizations navigating the challenges of WFA environments. Future work can build upon our findings by exploring larger, more diverse samples and analyzing long-term cybersecurity behaviour in remote settings. Further research could also address the privacy paradox, exploring the factors contributing to the discrepancy between employees’ concerns about cybersecurity and their actual behaviours.